OpenStudy (anonymous):

Jeremiah is factoring 0.64t^24 - 9 by using the rule, a^2 - b^2 = (a+B)(a-b). What will he use for the value of a? A.0.08t^12 B. 0.8t^12 C. 0.64 D. 0.8 I think it's B or C... Help please!

OpenStudy (sebastian_michaelis):

So basically 0.64t^24 – 9=a^2-b^2=(a+b)(a-b) Which number squared gives 0.64t^24 Compare and get your answer.

OpenStudy (anonymous):

Wait so it is b?

OpenStudy (sebastian_michaelis):

its C

OpenStudy (anonymous):

Ohh! Okay! Thanks for the help

OpenStudy (anonymous):

I got it wrong. The correct answer was B.